# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Three Independent Astrophysical Determinations of the Hubble Constant Using the James Webb Space Telescope

### v1

@@ 9.3 The Hubble Diagram
In Figure 10 we show the Hubble diagram for the CSP sample of 287 SNe Ia (blue filled circles ( CSP-I) and dark cyan filled circles ( CSP-II) where z0.01 and excluding super-Chandra (IA-SC) and IaX subtypes. In addition, two additional 3- outliers (SN2014D and SN2013hh) were excluded from the analysis. The calibrating galaxies provide distances only; their velocities are not used in the calibration of . Residuals from the fit in distance modulus are shown in the lower panel.
@@ 9.4 Markov Chain Monte Carlo Analysis
We explore the use of two Python packages, emcee (Foreman-Mackey et al., 2013), and pymc3 (Oriol et al., 2023) and find that they give consistent results. Initially, using emcee, a set of broad, uniform priors were adopted for 8 parameters (P0, P1, P2, , , Vpec, , and ) and 100 walkers were used. The burn-in time was set to be 5 times the autocorrelation time, , generally 500 steps, and the acceptance fraction was generally 0.45-0.49, within the range of 0.2-0.5 recommended by Gelman et al. (1996). The number of chains after burn-in was taken to be greater than 100, and convergence was visually checked in walker trajectory plots. The adopted parameter fits are given by the marginalized distributions and uncertainties quoted are 16th, 50th and 84th percentiles. In the case of the second analysis using pymc, 30,000 steps were implemented, and the burn-in time was set to 3,000. pymc uses “forward sampling” using the derivatives of the probability hyper-surface. Consistent results were obtained in both cases.
@@
We note that the sample of ten galaxies for which we have JWST data/distances for is currently small, resulting in a significant statistical uncertainty relative to longer-running HST-based studies (e.g., F19, R22). The scatter in the peak SN Ia magnitudes for this nearby sample is smaller than for the distant CSP sample. Allowing only for a single parameter to account for the scatter, , leads to an anomalously small dispersion in the P0 posterior, and hence, the posterior. We thus included an additional 9th parameter (), allowing for this difference in the scatter between the calibrating and distant sample, and thereby allowing for a larger uncertainty in . The scatter in P0 then increased by a factor of five, and the resulting value of increased by 1%, with a two-fold increase in its uncertainty. At present, it is not clear if the larger scatter for the more distant sample is a result of larger (unaccounted-for) errors (for example, errors in k-corrections, which are currently not included). Alternatively, the nearby sample may not be large enough to sample the true scatter. This situation will become clarified as further JWST data become available.
In Table 3 we list the parameter fits individually for the pymc runs for the Cepheid, TRGB and JAGB distances. We show an example corner plot from pymc for the TRGB calibrators in Figure 13. From our adopted pymc analysis, based on the -band CSP SN Ia data, the peak of the marginalized distributions for the three individual distance indicators are: = 72.05 1.86 for Cepheids, = 69.85 1.75 for the TRGB, and = 67.96 1.57 km s-1 Mpc-1 for the JAGB (statistical errors). These values are based on 10 SN Ia calibrators for the TRGB and Cepheids, and 7 calibrators for the JAGB method. The estimated statistical uncertainty calculated for the JAGB method (and listed in Table 3) is lower than that for the Cepheids and the TRGB simply as a consequence of the fact that the blinded sample of seven calibrating galaxies does not include three additional supernovae, SN 2011by, SN 2007sr, and SN 2012cg in the galaxies NGC 3972, NGC 4038 and NGC 4424, respectively. As described in §7.2, the radial measurements of the JAGB magnitude did not converge in these galaxies. In order not to give undue weight to the JAGB method, we increased its formal uncertainty (1.57) by the ratio of the average values (based on the TRGB and Cepheids) to that of the lower value for the JAGB: i.e., 0.157 0.13/0.11 = 1.85. A Bayesian analysis combining the three methods yields a value of = 69.96 1.05 km s-1 Mpc-1 (statistical error). We discuss the systematic uncertainty in §11.
Parameter | TRGB | JAGB | Cepheids |
P0 (mag) | -19.20 0.05 | -19.26 0.05 | -19.13 0.05 |
P1 (mag) | -0.90 0.10 | -0.90 0.10 | -0.90 0.10 |
P2 (mag) | -0.28 0.29 | -0.29 0.29 | -0.29 0.28 |
2.91 0.09 | 2.92 0.09 | 2.92 0.09 | |
(mag/dex) | 0.00 0.01 | 0.00 0.01 | 0.00 0.01 |
(mag) | 0.14 0.05 | 0.11 0.05 | 0.12 0.06 |
(mag) | 0.19 0.01 | 0.19 0.01 | 0.19 0.01 |
Vpec (km s-1) | 174 99 | 175 99 | 175 99 |
@@
H0 | 69.85 1.75 | 67.96 1.57aaAs described in the text, when combining the methods for the final analysis, the uncertainty for the JAGB method was increased by a factor of 1.18 (given by the ratio of the average values for TRGB and Cepheids to that of the lower value for the JAGB), so that the JAGB method does not unduly contribute more weight than the other two methods. | 72.05 1.86 |
@@
As previously noted, the total CSP sample includes a significant fraction of low-z SNe Ia which, if included in the MCMC analysis, would require a correction for the local density field. For example, based on predictions from the 2M++ density field of Carrick et al. (2015), Uddin et al. (2024) corrected the CSP peculiar velocity measurements (for which the average peculiar velocity is 440 km s-1). They found a mean sample velocity correction for the CSP survey of 90 km s-1, resulting in a net increase in of 0.55 km s-1 Mpc-1 in the B band. However, as shown in Brout et al. (2022), not including supernovae at z 0.01 avoids the bias due to nearby peculiar velocities. Also as previously noted, our preferred value of is based on a sample of 287 SNe Ia with z0.01, eliminating also super-Chandra (IA-SC) and IaX subtypes, thus requiring no additional correction to .
@@
In Table 4 we summarize the values of obtained using the three methods applied in this paper, based on the CSP SN Ia data. We show also values that are based on the Pantheon+ SN Ia data (see discussion in §9.5 below), for which there is good agreement. We present the results from different MCMC analyses (pymc and emcee), show the results adopting different redshift cuts for the CSP SNe Ia sample, adopting different nearby calibrators, calibrating the -band data for the CSP SNe Ia, and, for the Cepheids, with and without metallicity corrections. To within the uncertainties, the various cuts do not lead to significant variations in . The largest differences remain between the Cepheid compared to the TRGB and JAGB calibrations. We note that the total CSP sample of SNe Ia with -band data is not as large as that for the band (213 compared to 322), and the values of obtained based on the -band are larger than for the band. Uddin et al. (2024) searched, but found no evidence, for potential calibration errors. There are, however, hints of the existence of a brighter sub-population of SNe Ia in the near-infrared, which warrants further investigation. Finally, we list the combined, adopted result. The uncertainties in these values are discussed in §11.1.
Our adopted value of , based on the more extensive -band data is listed at the end of the table. We show the relative probability density functions (PDFs) for the three methods and the combined, adopted value in Figure 11. A Frequentist cumulative sum of the PDFs yields = 69.82 1.69 (stat) 1.12 (sys). A Bayesian analysis based on the product of the PDFs, adopting a flat prior, yields = 69.96 1.05 (stat) 1.12 (sys). A discussion of the uncertainties adopted follows in Table §5. We consider below the combined value of based on the total errors, but our preference for our adopted value is to keep the statistical and systematic errors separate, and easily identified.
In Figure 12 we plot the relative probability density functions for the three methods taking into account both the statistical and systematic uncertainties. In this case, the width of each Gaussian is given by . Assuming a flat prior, the Bayesian product of the PDFs yields = 69.59 1.58 (2.3%) km s-1 Mpc-1, based on all three methods (black curve). The JAGB and TRGB methods alone yield = 69.03 1.75 (2.5%) km s-1 Mpc-1 (blue-gray curve). We note that the latter is based solely on JWST data: that is, the calibration through NGC 4258 to the target galaxies is carried out with the same NIRCam detector. The Cepheid analysis relies on both HST and JWST data.
aafootnotemark: | Erroraafootnotemark: | Descriptionb,cb,cfootnotemark: | MCMC code |
@@
68.63 | +1.38 -1.32 | TRGB + CSP B band ; 9 SN calibratorsbbfootnotemark: , z0.01 | pymc |
68.62 | +1.41 -1.33 | TRGB + CSP B band ; 9 SN calibrators, z0.0065ccfootnotemark: | pymc |
68.72 | +1.38 -1.30 | TRGB + CSP B band ; 9 SN calibrators, z0ccfootnotemark: | pymc |
69.85 | +1.80 -1.70 | TRGB + CSP B band ; 10 SN calibrators, z0.01 | pymc |
68.03 | +1.48 -1.34 | TRGB + CSP B band ; 7 SN calibrators, z0.01 | pymc |
69.74 | +1.77 -1.71 | TRGB + CSP B band ; 10 SN calibrators, each with equal weight, z0.01 | pymc |
68.46 | +1.45 -1.34 | TRGB + CSP B band ; 9 SN calibrators, z0.01 | emcee |
69.64 | +1.73 -1.66 | TRGB + CSP B band ; 10 SN calibrators, z0.01 | emcee |
69.88 | +1.78 -1.74 | TRGB + CSP B band ; 10 SN calibrators, z0ccfootnotemark: | emcee |
@@
68.76 | +1.86 -1.76 | JAGB + CSP B band ; 9 SN calibratorsddfootnotemark: , z0.01 | pymc |
68.92 | +1.64 -1.58 | JAGB + CSP B band ; 10 calibrators, z0.01 | pymc |
67.96 | +1.60 -1.53 | JAGB + CSP B band ; 7 SN calibrators, z0.01 | pymc |
Cepheids (z0.01; CSP B band ) | |||
70.77 | 1.55 -1.55 | Cepheids (with metallicity correction) ; 9 SN calibrators | pymc |
70.76 | 1.55 -1.56 | Cepheids (with metallicity correction) ; 9 SN calibrators, z0.0065ccfootnotemark: | pymc |
70.82 | 1.57 -1.56 | Cepheids (with metallicity correction) ; 9 SN calibrators;, z0ccfootnotemark: | pymc |
72.05 | 1.85 -1.86 | Cepheids (with metallicity correction) ; 10 SN calibrators | pymc |
72.56 | 2.00 -2.03 | Cepheids (without metallicity correction) ; 10 SN calibrators | pymc |
71.35 | +1.50 -1.58 | Cepheids (without metallicity correction) ; 9 SN calibrators | pymc |
Combined Results (z0.01, CSP B-band) | |||
69.82 | 1.69 1.12 | Combined (Cepheids, TRGB, JAGB) ; 10, 10, 8 SN calibrators; Frequentist | pymc |
69.96 | 1.05 1.12 | Combined (Cepheids, TRGB, JAGB) ; 10, 10, 7 SN calibrators; Bayesian | pymc |
69.03 | 1.75 | Combined (TRGB, JAGB alone) ; 11, 7 SN calibrators; Bayesian | pymc |
@@
69.94 | … | TRGB + Pantheon+ B band ; 11 SN calibrators | SHoES eefootnotemark: |
68.56 | … | JAGB + Pantheon+ B band ; 8 SN calibrators | SHoES eefootnotemark: |
71.95 | … | Cepheids + Pantheon+ B band ; 11 SN calibrators | SHoES eefootnotemark: |
H-band Results | |||
74.71 | +1.34 -1.35 | Cepheids CSP H band; 10 SN calibrators, z 0.01 | |
71.59 | +1.85 -1.76 | TRGB + CSP H band ; 10 SN calibrators, z 0.01 | |
70.96 | +1.91 -1.81 | TRGB + CSP H band ; 9 SN calibrators, z 0.01 | |
69.34 | +1.71 -1.62 | JAGB + CSP H band ; 7 SN calibrators, z 0.01 | |
@@
69.96 1.05 (stat) 1.12 (sys) |
In summary, the final value that we adopt for this paper is based on all three calibrators, tying in to the -band CSP SN Ia sample. As noted, there is a systematic offset between the Cepheid distance scale and the JAGB/TRGB distance scale. We have listed separately the values obtained for the combination of TRGB and JAGB calibrations, and that from Cepheids. It will be important for future studies to resolve this local distance-scale discrepancy.
@@ 9.5 Comparison with Pantheon+
In the Pantheon+ analysis, Scolnic et al. (2022) collected and cross-calibrated the data for 1550 individual SNe Ia, superseding earlier Pantheon (Scolnic et al., 2018) and Joint Light-Curve (Betoule et al., 2014) analyses. The total sample includes SNe Ia in the redshift range 0 z 2.3. The analysis aims to standardize the B-band photometry from 18 individual surveys obtained with a wide variety of telescopes and instruments777 The Pantheon+ catalog is available at https://github.com/PantheonPlusSH0ES/DataRelease.. In the SHoES determination of , (e.g., Riess et al., 2009, 2022), a simultaneous fit is undertaken for the Cepheid and SNe Ia data, minimizing a statistic, and most recently providing the covariances. R22 also utilize emcee as a check of their methodology, finding almost exact agreement. The SHoES Cepheid calibration of the Pantheon+ SNe Ia sample from Scolnic et al. (2022) results in a value of = 73.04 1.04 km s-1 Mpc-1 for the 277 SNe Ia with 0.023 z 0.15.
In Figure 14 we compare the absolute magnitude SN Ia peak brightness, MB, for the set of SNe Ia contained in the Pantheon+ sample for which there are R22 Cepheid distances, and that now have distances also measured based on the three methods from this study. There are 10 galaxies for which we have measured TRGB and Cepheid distances (hosting 11 SNe Ia), and 7 galaxies with well-measured JAGB distances (hosting 8 SNe Ia). The weighted mean magnitudes and error on the mean for the three samples are = -19.286 0.029 mag (Cepheids), -19.345 0.026 mag (TRGB), and -19.384 0.023 mag (JAGB), respectively. These values can be compared to the overlapping sample of R22 SN Ia discussed above where, = -19.315 0.032 mag.
@@
For comparison purposes, to be completely consistent with the SN Ia analyses undertaken by Scolnic et al. (2022) and R22, we do not re-analyze the Pantheon+ data. We determine by scaling the absolute MB values relative to those measured by SHoES, allowing for the fact that our calibration is based on NGC 4258 alone, and not the baseline calibration of the LMC, Milky Way and NGC 4258. The R22 value of based on NGC 4258 alone is 72.51 1.54 km s-1 Mpc-1 (which can be compared to their baseline value of 73.04 1.01 km s-1 Mpc-1). In the case of the Pantheon+ data we find = 71.95 Cepheids, = 69.94 for the TRGB, and = 68.56 km s-1 Mpc-1 for the JAGB. These values are in excellent agreement with those determined based on the CSP SNe Ia calibration.

### v2

@@ 10.3 The Hubble Diagram
In Figure 9 we show the Hubble diagram for the CSP sample of 287 SNe Ia (blue filled circles ( CSP-I) and dark cyan filled circles ( CSP-II) where z0.01, while excluding super-Chandra (IA-SC) and IaX subtypes. In addition, two additional 3 outliers (SN2014D and SN2013hh) were excluded from the analysis. The calibrating galaxies provide distances only; their velocities are not used in the calibration of . (That is, these galaxies would appear in Figure 9 to fall on the trend line with zero dispersion. The calibrating SNe Ia constrain only the intercept.) Residuals from the fit in distance modulus are shown in the lower panel.
@@ 10.4 Markov Chain Monte Carlo Analysis
We explore the use of two Python packages, emcee (Foreman-Mackey et al., 2013) and pymc3 (Oriol et al., 2023), and find that they give consistent results. Initially, using emcee, a set of broad, uniform priors were adopted for 8 parameters (P0, P1, P2, , , Vpec, , and ) and 100 walkers were used. The burn-in time was set to be 5 times the autocorrelation time, , generally 500 steps, and the acceptance fraction was generally 0.45-0.49, within the range of 0.2-0.5 recommended by Gelman et al. (1996). The number of chains after burn-in was taken to be greater than 100, and convergence was visually checked in walker trajectory plots. The adopted parameter fits are given by the marginalized distributions and uncertainties quoted are 16th, 50th and 84th percentiles. In the case of the second analysis using pymc, 30,000 steps were implemented, and the burn-in time was set to 3,000. pymc uses “forward sampling” using the derivatives of the probability hyper-surface. Consistent results were obtained in both cases.
@@
We note that the sample of ten galaxies for which we have JWST distances (and 11 SNe Ia) is currently small, resulting in a larger statistical uncertainty relative to longer-running HST-based studies (e.g., F19, R22). The scatter in the peak SN Ia magnitudes for this nearby sample is smaller than for the distant CSP sample. Allowing only for a single parameter to account for the scatter, , leads to an anomalously small dispersion in the P0 posterior, and hence, the posterior. We thus included an additional 8th parameter (), allowing for this difference in the scatter between the calibrating and distant sample, and thereby allowing for a larger uncertainty in . The scatter in P0 then increased by a factor of five, and the resulting value of increased by 1%, with a two-fold increase in its uncertainty.
Table 4 presents the parameter fits derived from our pymc analysis, encompassing results for JWST TRGB, JAGB, and combined TRGB+JAGB distances, as well as the full HST +JWST TRGB sample. Using B-band CSP SN Ia data and the new JWST data, we obtained the following peak marginalized distributions: 68.81 1.79 km s-1 Mpc-1 for the TRGB and 67.80 1.49 km s-1 Mpc-1 for JAGB (statistical errors). These values are based on 11 SN Ia calibrators for TRGB and 8 for the JAGB. In an earlier version of this paper, the analysis did not include SN 2017cbv, which was observed after the CSP-II program had been completed. SN 2017cbv was, however, observed with exactly the same instrumentation and optical bandpasses as for CSP-II (see Burns et al., 2020).
The statistical uncertainty for the JAGB method, as listed in Table 4, is smaller than that for TRGB. This is due to the blinded JAGB sample excluding three supernovae (SN 2011by, SN 2007sr, and SN 2012cg) that are included in the TRGB sample. These omissions, resulting from radial measurements not converging for these galaxies, led to an artificially small dispersion in the JAGB-SN Ia calibration. To avoid underestimating the JAGB uncertainty, we increased it by a factor of 2.45, the ratio of the calibration dispersions between TRGB and JAGB.
A pymc analysis of the weighted mean of JWST TRGB and JAGB distances (Table 2, column 7) yields an value of 69.02 1.66 km s-1 Mpc-1 (statistical error). The differences in between TRGB, JAGB, and the combined sample arise from their respective uncertainties and weightings. Notably, the averaged distance moduli for NGC 5643 and NGC 2442 are significantly lower in the combined sample than for the TRGB alone, contributing to a slight (0.3%) increase in .
Expanding the analysis to the larger combined HST + JWST TRGB sample of 24 SNe Ia provides our best current estimate, yielding = 70.39 1.22 km s-1 Mpc-1 (statistical error). We show an example corner plot from pymc for the 24 TRGB calibrators in Figure 10.
The values obtained from the TRGB sample (68.81 1.79) and the full HST+JWST sample (70.39 1.22) are statistically consistent. Eleven of the 24 supernovae are common to both samples, and accounting for uncorrelated errors in the independent measurements results in a difference at the 0.8 level ((70.39 - 68.81) / ).
A discussion of the systematic uncertainties is deferred to a later section of the paper. We keep the statistical and systematic errors separate, so that they can be easily identified.
| Parameter | TRGBtot | TRGBJWST | JAGBJWST | TRGB+JAGB |
| No. of calibrators | 24 | 11 | 8 | 11 |
| P0 (mag) | -19.18 0.04 | -19.23 0.05 | -19.26 0.05 | -19.22 0.05 |
| P1 (mag) | -0.93 0.10 | -0.91 0.10 | -0.91 0.10 | -0.91 0.10 |
| P2 (mag) | -0.32 0.28 | -0.32 0.29 | -0.32 0.29 | -0.32 0.29 |
| 2.90 0.09 | 2.92 0.09 | 2.92 0.09 | 2.92 0.09 | |
| (mag/dex) | 0.00 0.01 | 0.00 0.01 | 0.01 0.01 | 0.01 0.01 |
| (mag) | 0.15 0.03 | 0.16 0.05 | 0.11 0.05 | 0.15 0.05 |
| (mag) | 0.19 0.01 | 0.19 0.01 | 0.19 0.01 | 0.19 0.01 |
@@
| H0 ( km s-1 Mpc-1) | 70.39 1.22 | 68.81 1.80 | 67.80 2.17aaAs described in the text, the uncertainty for the JAGB method was increased by a factor of 1.45 (given by the ratio of the value for TRGB to that of the lower value for the JAGB), so that the JAGB method does not unduly contribute more weight simply because of the smaller dispersion in its smaller number of calibrators. | 69.02 1.56 |
@@
As previously noted, the total CSP sample includes a significant fraction of low-redshift SNe Ia which, if included in the MCMC analysis, would require a correction for the local-density-field velocity perturbations. For example, based on predictions from the 2M++ density field of Carrick et al. (2015), Uddin et al. (2024) corrected the CSP peculiar velocity measurements (for which the average peculiar velocity is 440 km s-1). They found a mean sample velocity correction for the CSP survey of 90 km s-1, resulting in a net increase in of 0.55 km s-1 Mpc-1 in the B band. However, as shown in Brout et al. (2022), not including supernovae at z 0.01 avoids the bias due to nearby peculiar velocities. Also as previously noted, our preferred value of is based on a sample of 287 SNe Ia with z0.01, eliminating also super-Chandra (IA-SC) and IaX subtypes, thus requiring no additional correction to .
@@
In Table 5 we summarize the values of obtained using the two methods applied in this paper, based on the CSP SN Ia data. We show also values that are based on the Pantheon+ SN Ia data (see discussion in §10.5 below), for which there is good agreement. We present the results from different MCMC analyses (pymc and emcee), show the results adopting different redshift cuts for the CSP SNe Ia sample, adopting different nearby calibrators (excluding those which show the largest deviations, for example), and for calibrating the -band data for the CSP SNe Ia. We have also examined the difference in when excluding SNe Ia observed as part of the CSP, but not the SHoES analysis; i.e., SNe Ia that occurred in elliptical or S0 galaxies. To within the uncertainties, the various cuts do not lead to significant variations in , and all are consistent with a value of = 70 km s-1 Mpc-1. In no case do we find a value of of 73 or 74 km s-1 Mpc-1. We note that the total CSP sample of SNe Ia with -band data is not as large as that for the band (213 compared to 322), and the values of obtained based on the -band are larger than for the band. Uddin et al. (2024) searched, but found no evidence, for potential calibration errors. There are, however, hints of the existence of a brighter sub-population of SNe Ia in the near-infrared, which warrants further investigation. The uncertainties in these values are discussed in §13.4.
Our adopted value of , based on the more extensive -band data, and the combined (HST + JWST) TRGB sample is listed at the end of the table. We do not apply a peculiar-velocity model correction to this value of . Uncertainties in model-dependent corrections are difficult to quantify, and may be as large, or larger, than the corrections themselves.888For example, in the extensive fundamental plane study undertaken as part of the DESI Peculiar Velocity Survey, in their Figure 7, Said et al. (2024) plot several velocity models (e.g., Carrick et al., 2015) in comparison to the residuals from their Hubble diagram, none of which provide a good fit to the data. Uddin et al. (2024) apply a 0.55 km s-1 Mpc-1 correction to their value of ; here we absorb the uncertainty in the velocity modeling as an additional systematic uncertainty. To be conservative, we take this uncertainty, , to be an additional 1%, which also provides allowance for remaining uncertainty in the photometric calibration for the SNe Ia.
| aafootnotemark: | Erroraafootnotemark: | Description | MCMC code |
@@
| 70.39 | +1.23 -1.21 | TRGB + CSP B band ; 24 SN calibrators, z0.01 ; 291 distant SN | pymc |
| 70.39 | +1.23 -1.22 | TRGB + CSP B band ; 24 SN calibrators, z0.0065 ; 303 distant SN | pymc |
| 70.48 | +1.24 -1.22 | TRGB + CSP B band ; 24 SN calibrators, z0.015 ; 264 distant SN | pymc |
| 70.37 | +1.26 -1.25 | TRGB + CSP B band ; 24 SN calibrators, z0.02 ; 235 distant SN | pymc |
| 69.81 | +1.32 -1.29 | TRGB + CSP B band ; 24 SN calibrators, z0.025 ; 201 distant SN | pymc |
| 70.11 | +1.24 -1.23 | TRGB + CSP B band ; 23 SN calibrators, omit SN2007on, z0.01 | pymc |
| 70.01 | +1.21 -1.19 | TRGB + CSP B band ; 23 SN calibrators, omit SN2011by, z0.01 | pymc |
| 70.73 | +1.21 -1.18 | TRGB + CSP B band ; 23 SN calibrators, omit SN1981D z0.01 | pymc |
| 70.45 | +1.23 -1.19 | TRGB + CSP B band ; 22 SN calibrators, omit SN1981D and 2007on, z0.01 | pymc |
| 69.96 | +1.18 -1.16 | TRGB + CSP B band ; 23 SN calibrators, omit SN2007af, z0.01 | pymc |
| 70.50 | +1.54 -1.49 | TRGB + CSP B band ; 16 SN calibrators, omit SNe not observed by R22bbfootnotemark: , z0.01 | pymc |
| 68.81 | +1.80 -1.77 | TRGB + CSP B band ; 11 SN calibrators, JWST distances only, z0.01 | pymc |
| 70.43 | +1.22 -1.30 | TRGB + CSP B band ; 24 SN calibrators, z0.01 | emcee |
@@
| 68.75 | +1.56 -1.52 | JAGB + CSP B band ; 11 SN calibrators, z0.01 | pymc |
| 67.80 | +1.50 -1.48 | JAGB + CSP B band ; 8 SN calibratorsccfootnotemark: , z0.01 | pymc |
| TRGB + JAGB Distances Averaged (z0.01, CSP B-band) | |||
| 69.02 | 1.69 -1.63 | Averaged (TRGB, JAGB) distances ; 11 SN calibrators | pymc |
| 67.60 | 1.58 -1.51 | Averaged (TRGB, JAGB) distances ; 8 SN calibrators | pymc |
@@
| 69.29 | … | TRGB + Pantheon+ B band ; 11 SN calibrators | SHoES ddfootnotemark: |
| 68.80 | … | JAGB + Pantheon+ B band ; 8 SN calibrators | SHoES ddfootnotemark: |
@@
| 70.39 1.22 (stat) 1.33 (sys) 0.70 () | |||
In summary, the final value that we adopt for this paper is based on the HST + JWST TRGB sample of 24 SN Ia calibrators, tying into the -band CSP SN Ia sample with = 70.39 1.22 (stat) 1.33 (sys) 0.70 []. We note the good agreement of the values using only the JWST TRGB and JAGB distances, as well as their averaged distances, despite the lower statistical power of the JWST samples alone.
@@ 10.5 Comparison with Pantheon+
In the Pantheon+ analysis, Scolnic et al. (2022) collected and cross-calibrated the data for 1550 individual SNe Ia, superseding earlier Pantheon (Scolnic et al., 2018) and Joint Light-Curve (Betoule et al., 2014) analyses. The total sample includes SNe Ia in the redshift range 0 z 2.3. The analysis aims to standardize the B-band photometry from 18 individual surveys obtained with a wide variety of telescopes and instruments999 The Pantheon+ catalog is available at https://github.com/PantheonPlusSH0ES/DataRelease.. In the SHoES determination of , (e.g., Riess et al., 2009, 2022), a simultaneous fit is undertaken for the Cepheid and SNe Ia data, minimizing a statistic, and most recently providing the covariances. R22 also utilize emcee as a check of their methodology, finding almost exact agreement. The SHoES Cepheid calibration of the Pantheon+ SNe Ia sample from Scolnic et al. (2022) results in a value of = 73.04 1.04 km s-1 Mpc-1 for the 277 SNe Ia with 0.023 z 0.15.
@@
For comparison purposes, to be completely consistent with the SN Ia analyses undertaken by Scolnic et al. (2022) and R22, we do not re-analyze the Pantheon+ data. We determine by scaling the absolute MB values relative to those measured by R22 (their ’baseline’ fit) where MB = -19.253 0.027 mag, allowing for the fact that our calibration is based on NGC 4258 alone, and not the baseline calibration of the LMC, Milky Way and NGC 4258, which corresponds to value of = 73.04 1.01 km s-1 Mpc-1. The R22 value of based on NGC 4258 alone is 72.51 1.54 km s-1 Mpc-1. That is,
We note that this value is consistent with that of R22, Table 5, entry #10, but differs from that quoted by R24b who quotes instead MB = -19.28. Adopting a value of MB = -19.269 for the Pantheon+ data we find = 69.29 for the TRGB and = 68.80 km s-1 Mpc-1 for the JAGB. These values are in excellent agreement with those determined based on the CSP SNe Ia calibration.
10.6 Absolute Magnitude MB Distributions for Pantheon+ SN Ia Calibrators
In Figure 11 we compare the absolute magnitude SN Ia peak brightness, MB, for the set of SNe Ia contained in the Pantheon+ sample. There are 20 galaxies for which we have measured TRGB distances (hosting 24 SNe Ia), and 7 galaxies with well-measured JAGB distances (hosting 8 SNe Ia). The weighted mean magnitude and error on the mean for the total sample is = -19.316 0.033 and -19.339 0.029 mag (with and without SN 2007af, respectively). This single supernova has a disproportionate effect: including SN2007af increases by 1%; i.e., it moves the weighted average by nearly the entire size of the R22 error bar. These values can be compared to the overlapping sample of R22 SN Ia discussed above where = -19.317 0.033 mag. Finally, the weighted mean magnitudes and error on the mean are = -19.363 0.026 mag (TRGB), and -19.377 0.023 mag (JAGB), respectively.
We show in Figure 12 a histogram of absolute magnitude (MB) values for the 24 host-SNe Ia calibrating galaxies in the CCHP sample. The TRGB distances are those from this paper and F19, and the apparent SN Ia magnitudes (mB) values are those adopted by R22. Shown in yellow are the JWST data from this paper. The expanded HST + JWST TRGB sample of 24 SNe Ia now exhibits a distribution more representative of the larger R22 sample of 42 SNe Ia, consistent with the expected ‘reversion to the mean’ for larger datasets. This increased sample size demonstrates that the mean of the TRGB-calibrated MB sample has remained stable. The criticism from Riess et al. (2024b) [hereafter (R24)] regarding our earlier, smaller JWST-only sample is no longer applicable, as it is superseded by our current analysis of 24 SNe Ia using HST and JWST data.
For comparison, we show in Figure 13 a histogram of the absolute magnitude (MB) values for the host-SNe Ia calibrating galaxies in the SHoES sample. The Cepheid distances and apparent magnitudes (mB) values are again from R22. The first point to note from Figures 12 and 13 is their overall similarity. The range of MB values is comparable, and both exhibit central peaks. The two histograms have been plotted to the same scale. Shown in yellow are the 12 SNe Ia that have the highest weight in the calibration. These 12 SNe Ia alone contribute half the weight of the SHoES sample.
In Figure 14 we plot the MB histograms for the CSP distant sample of SNe Ia and overplot (for a value of = 70 km s-1 Mpc-1) the nearby TRGB calibrators, this time on the CSP magnitude scale defined in §10.2 above, and scaled by a factor of ten.
In terms of overall shape, this figure illustrates that the nearby TRGB SN Ia sample is a very reasonable match to the more distant, and larger SN Ia flow sample. A Kolmogorov-Smirnov (KS) test comparing the two distributions yields a p-value of 0.12, indicating that the two distributions are consistent in shape. The same conclusion holds if we normalize the two histograms to the same area, and shift the nearby calibration sample by 0.03 mag (or 1% in the SHoES value of ). Thus, we conclude that the TRGB sample is not biased (or skewed) with respect to the more distant sample, and it can provide a reliable calibration for the determination of .